Determine the [OH−] , pH, and pOH of a solution with a [H+] of 4.0×10−11 M at 25 °C.
1
Expert's answer
2020-04-15T01:52:10-0400

Kw = [H+][OH-] = 1*10-14

"pH = -lg[H^+]=-log_{10}(4*10^{-11})=10.4"

"pOH = 14-pH = 3.6"

"[OH^-]=10^{-pOH}=0.00025=2.5*10^{-4}"